# QOS: A Quantum Operating System

### v1

@@ 9.3. Estimator
Estimator’s Accuracy. Figure 12 (c) shows the fidelity of the eight benchmarks when run on QPUs selected by the estimator versus when run on the IBM Auckland QPU. The QPU selected for the BV benchmark is Auckland; therefore, we omit this result. For the rest of the benchmarks, the IBM Sherbrooke and Brisbane QPUs were automatically selected. Interestingly, the fidelity is on par or even higher than IBM Auckland, except for only one benchmark, the QAOA-P1.
RQ2 takeaway: QOS’s estimator automatically identifies QPUs with higher fidelity than the current standard practice.
9.4. Multi-programmer
@@
Utilization vs. Fidelity. Figure 13 (a) shows the average fidelity of nine benchmarks with utilization of 30%, 60%, and 88%. The three bars represent: no multi-programming (No M/P) refers to large circuits that run solo, baseline multi-programming (Baseline M/P) refers to (Das et al., 2019), and QOS’s multiprogramming approach (QOS M/P). There is an average 9.6$\times$ improvement in fidelity compared to solo execution and an average 15% (1.15$\times$) improvement compared to the baseline.
Figure 13. QOS Multi-programmer (§ 9.4). (a) Impact of multi-programming on fidelity. There is a 9.6$\times$ and 1.15$\times$ improvement compared to no multi-programming and the baseline, respectively. (b) Effective utilization. There is 7.2% higher effective utilization on average. (c) Relative fidelity w.r.t. solo circuit execution. There is an average 9.6% drop in fidelity due to QOS’s multi-programming.
Effective Utilization. The results in Figure 13 (b) show that QOS achieves, on average, a 7.2% higher effective utilization (§ 8.2), with a maximum improvement of 10.1%.
Fidelity Penalty vs. Solo Execution. In Figure 13 (c), we evaluate the fidelity penalty of multiprogramming vs. solo circuit execution for utilization of 30%, 60%, and 88%. The fidelity loss is 2%, 9%, and 18%, respectively. The average fidelity loss is 9.6% compared to solo execution, which is in line with previous studies (Das et al., 2019; Liu and Dou, 2021). In the worst case (18%), the fidelity loss is caused by the restrictions in high-quality qubit allocations and the crosstalk errors.
Figure 14. QOS Scheduler (§ 9.5). (a) Formula-based scheduling policy: Average fidelity vs. average waiting times. A fidelity weight $c=0.7$ achieves 5$\sim$ 5$\times$ 5 × lower waiting time for only $\sim 2\%$ lower fidelity. (b) Genetic algorithm policy: it creates a Pareto front of schedules, where a fidelity weight $c=0.5$ achieves 2×2$\times$2 × lower waiting times for $\sim 4\%$ fidelity decrease. (c) QPU load as the total runtime of each QPU for the formula-based policy. The maximum load difference between any two QPUs is $15.2\%$.
RQ3 takeaway: The QOS multi-programmer improves fidelity by 1.15–9.6$\times$ and effective utilization by 7.2% compared to the baselines while incurring an acceptable fidelity penalty ($<10\%$) compared to solo execution.
9.5. Scheduler
@@
Workload Generation. To generate realistic workloads, we monitored all available QPUs on the IBM Quantum Cloud (ibmQuantum, [n. d.]) for ten days in November 2023 to estimate the hourly job arrival rate. The average hourly rate is 1500 jobs per hour and is the baseline system workload for our evaluation.
@@
RQ4 takeaway: QOS scheduler balances the trade-off between waiting times and fidelity by reducing them 5$\times$ and only 2%, respectively, while balancing the load across QPUs.
10. Related work
Quantum optimization techniques can be categorized as (1) qubit mapping and routing (Li et al., 2019; Murali et al., 2019; Tannu and Qureshi, 2019c; Zhang et al., 2021; Molavi et al., 2022; Tannu and Qureshi, 2019a; Zulehner et al., 2019; Wille et al., 2019; Siraichi et al., 2018; Patel et al., 2022a; Patel and Tiwari, 2021; Baker et al., 2021), (2) instruction/pulse scheduling (Das et al., 2021b; Tripathi et al., 2022; Zlokapa and Gheorghiu, 2020; Smith et al., 2022; Murali et al., 2020; Shi et al., 2019a; Gokhale et al., 2020), (3) gate synthesis/decomposition (Shi et al., 2019a; Patel et al., 2022a; Litteken et al., 2023; Das et al., 2023; Xu et al., 2023; Patel et al., 2022b; Shi et al., 2019b), (4) execution post-processing and readout improvement (Bravyi et al., 2021; Maciejewski et al., 2020; Patel and Tiwari, 2020b; Das et al., 2021a; Patel and Tiwari, 2020a; Dangwal et al., 2023; Tannu and Qureshi, 2019b), and (5) circuit compaction (Mitarai and Fujii, 2021; Peng et al., 2020; Tang et al., 2021b; Bello et al., 2023; Ayanzadeh et al., 2023; Bichsel et al., 2020; Paradis et al., 2021; Ding et al., 2020; Paler et al., 2016; Hua et al., 2023). These techniques are implemented standalone without a compiler infrastructure and, thus, are not composable. Instead, the QOS compiler offers a powerful IR that enables incorporating such techniques in a composable manner.
Moreover, application-specific optimizations focus only on specific algorithms to enhance fidelity but lack generality (Alam et al., 2020; Lao and Browne, 2022; Gokhale et al., 2019b, a; Li et al., 2022; Stein et al., 2022; Resch et al., 2021; Hao et al., 2023; Tüysüz et al., 2023; Ravi et al., 2022a, 2023). In contrast, the QOS compiler is a generic approach applicable to all applications.
@@ 10. Related work
The state-of-the-art in quantum multi-programming is fairly limited (Das et al., 2019; Liu and Dou, 2021) and focuses solely on high-quality mapping, overlooking the systematic selection of compatible programs for utilization or fidelity. Notably, key optimizations from (Das et al., 2019) are integrated into the Qiskit transpiler workflow (qiskit-transpiler, [n. d.]), therefore, are already used by the QOS Virtualizer (§ 7.2).
Lastly, current quantum scheduling methods (Ravi et al., 2021b; Salm et al., 2022; Bhoumik et al., 2023) are limited because they (1) schedule circuits one at a time, (2) neglect QPU utilization, (3) lack fine control over waiting times versus fidelity, and (4) require manual input for final scheduling decisions. Work in the quantum cloud computing area (Karalekas et al., 2020; Leymann et al., 2020; Ravi et al., 2022b) and in quantum serverless (Nguyen et al., 2022; Garcia-Alonso et al., 2022; Grossi et al., 2021); describes quantum cloud characteristics or potential architectures, but QOS is the first end-to-end QPU management system.
11. Conclusion
We presented QOS, a system that composes cross-stack OS abstractions to address the challenges of quantum computing holistically. The synergy between compaction techniques, performance estimation, multi-programming, and scheduling systematically explores the tradeoff space associated with quantum. Specifically, QOS achieves up to 456.5×456.5$\times$456.5 × higher fidelity at a 12×12$\times$12 × overhead cost, up to 9.6×9.6$\times$9.6 × higher fidelity for a target utilization for $9.6\%$ lower fidelity than solo execution, and up to 5×5$\times$5 × lower waiting times for $2\%$ lower fidelity.
Contributions. Our main contributions include:
  1. (1)
    To our knowledge, QOS is the first attempt to combine circuit compaction with quantum resource management to tackle the challenges of QPUs holistically.
  2. (2)
    We leverage the QOS compiler infrastructure to compose optimizations that improve fidelity in a scalable manner, significantly outperforming their individual application (i.e., the current practice).
  3. (3)
    To our knowledge, we are the first to account for and improve both temporal and spatial QPU utilization when multi-programming quantum programs, while mitigating its associated fidelity penalties.
  4. (4)
    Our scheduler balances the inherent tradeoff between fidelity and waiting times, leading to better overall QoS.
@@ Acknowledgements
We thank Karl Jansen and Stefan Kühn from the Center for Quantum Technology and Applications (CQTA)- Zeuthen for supporting this work by providing access to IBM quantum resources. We also thank Ahmed Darwish and Dmitry Lugovoy for their contributions to this work. Funded by the Bavarian State Ministry of Science and the Arts as part of the Munich Quantum Valley (MQV).

### v2

@@ 9.3 Estimator
Estimator’s accuracy. Figure 10 (c) shows the fidelity of the eight benchmarks when run on QPUs selected by the estimator versus when run on the IBM Auckland QPU. The QPU selected for the BV benchmark is Auckland; therefore, we omit this result. For the rest of the benchmarks, the IBM Sherbrooke and Brisbane QPUs were automatically selected. Interestingly, the fidelity is on par or even higher than IBM Auckland, except for only one benchmark, the QAOA-P1, possibly because of its unique, power-law connectivity structure.
RQ2 takeaway. QOS’s estimator automatically identifies QPUs with higher fidelity than the current standard practice.
9.4 Multi-programmer
@@
Figure 11: Multi-programmer (§ 9.4). (a) Impact of multi-programming on fidelity. There is a 9.6$\times$ and 1.15$\times$ improvement compared to no multi-programming and the baseline, respectively. (b) Effective utilization. There is 7.2% higher effective utilization on average. (c) Relative fidelity w.r.t. solo circuit execution. There is an average 9.6% drop in fidelity due to QOS’s multi-programming.
Utilization vs. fidelity. Figure 11 (a) shows the average fidelity of nine benchmarks with utilization of 30%, 60%, and 88%. The three bars represent: no multi-programming (No M/P) refers to large circuits that run solo, baseline multi-programming (Baseline M/P) refers to [16], and QOS’s multi-programming approach (QOS M/P). There is an average 9.6$\times$ improvement in fidelity compared to solo execution and an average 15% (1.15$\times$) improvement compared to the baseline.
Effective utilization. The results in Figure 11 (b) show that QOS achieves, on average, a 7.2% higher effective utilization (§ 7), with a maximum improvement of 10.1%. We attribute this improvement to the inclusion of temporal utilization.
Fidelity penalty vs. solo execution. In Figure 11 (c), we evaluate the fidelity penalty of multiprogramming vs. solo circuit execution for utilization of 30%, 60%, and 88%. The fidelity loss is 2%, 9%, and 18%, respectively. The average fidelity loss is 9.6% compared to solo execution, which is in line with previous studies [16, 40]. In the worst case (18%), the fidelity loss is caused by the restrictions in high-quality qubit allocations and the crosstalk errors.
Figure 12: Scheduler (§ 9.5). (a) Formula-based scheduling policy: Average fidelity vs. average waiting times. A fidelity weight $c=0.7$ achieves 5$\sim$ 5$\times$ 5 × lower waiting time for only $\sim 2\%$ lower fidelity. (b) Genetic algorithm policy: it creates a Pareto front of schedules, where a fidelity weight $c=0.5$ achieves 2×2$\times$2 × lower waiting times for $\sim 4\%$ fidelity decrease. (c) QPU load as the total runtime of each QPU for the formula-based policy. The maximum load difference between any two QPUs is $15.2\%$.
RQ3 takeaway. The QOS multi-programmer improves fidelity by 1.15–9.6$\times$ and effective utilization by 7.2% compared to the baselines while incurring an acceptable fidelity penalty ($<10\%$) compared to solo execution.
9.5 Scheduler
@@
Workload generation. To generate realistic workloads, we monitored all available QPUs on the IBM Quantum Cloud [32] for ten days in November 2023 to estimate the hourly job arrival rate. The average hourly rate is 1500 jobs per hour and is the baseline system workload for our evaluation.
@@
RQ4 takeaway. QOS scheduler balances the trade-off between waiting times and fidelity by reducing them 5$\times$ and only 2%, respectively while balancing the load across QPUs.
10 Related work
Quantum error mitigation. Error mitigation techniques can be categorized as (1) execution pre- and post-processing [52, 15, 51, 13, 75, 74, 14, 53, 71, 54, 42], (2) circuit divide-and-conquer [73, 6, 4, 54], and (3) qubit reuse [29, 48, 8, 49, 48, 34]. Unfortunately, all these techniques are implemented standalone without any infrastructure to compose them. Instead, QOS integrates and composes at least one technique per category in a single software stack, achieving higher fidelity and abstracting away the complexity from the programmer.
Quantum performance estimation. Performance estimation in the context of quantum computing exhibits preliminary work, focusing specifically on quantum resource estimation [7, 59] and fidelity estimation [47]. The former work estimates the number of physical qubits required to run an input job given pre-defined fidelity and error correction assumptions. However, QOS focuses on the opposite task: given the hardware constraints, it estimates the fidelity of the input quantum job. The latter work has a similar focus and utilizes similar techniques; however, it does not support ML-based approaches or non-analytical-model methods.
@@ 10 Related work
Quantum multi-programming. Quantum multi-programming work [16, 40] focuses solely on mapping two circuits on a single QPU while providing similar fidelity (allocation fairness). Both works overlook the systematic selection of compatible programs for utilization or fidelity and do not account for temporal utilization. Additionally, key optimizations from [16] are integrated into the Qiskit transpiler workflow [58], and therefore, are already used by QOS.
Quantum job scheduling. Current quantum scheduling methods [62, 64, 72, 81] are limited because they (1) only perform single-to-many scheduling, (2) do not account for QPU utilization, (3) lack fine control over the fundamental design tradeoffs (§ 3.4), or (4) require manual input for final scheduling decisions. In contrast, QOS multiplexes circuits across space and time in a many-to-many fashion, increases QPU utilization, balances the inherent tradeoffs of quantum, and abstracts away the underlying resource management.
Cloud OSes, resource management, and job scheduling. Classical cloud operating, resource management, and job scheduling systems have been extensively researched in the past decades [28, 80, 66, 36, 79, 84, 67, 83, 56, 46, 37]. However, the classical domain, even when addressing accelerator heterogeneity, does not face the unique challenges of quantum computing (§ 3) and, as such, cannot be trivially adapted to accommodate its needs.
11 Conclusion and Discussion
We presented QOS, the first quantum operating system to (1) holistically address quantum computing challenges through a modular, policy-mechanism-separated architecture, (2) enable cross- and intra-layer optimizations via end-to-end co-design, (3) systematically explore key tradeoffs between fidelity, utilization, and waiting times, and (4) introduce novel abstractions and techniques across the stack, including the Qernel, composable error mitigation, compatibility-based multi-programming, and multi-objective scheduling.
We believe that QOS lays the foundation for a quantum operating system whose core principles are designed to persist even as hardware matures. While advances in QPU technology will improve qubit counts and reliability, fundamental challenges—such as noise, heterogeneity, and resource scarcity—will continue to demand careful tradeoffs between fidelity, utilization, and waiting times. Looking forward, QOS’s principles extend naturally to fault-tolerant quantum computing (FTQC): QOS’s fidelity-aware scheduler and compatibility-driven multi-programmer can be adapted to logical qubit placement, routing, and error-corrected resource allocation, preserving execution efficiency while accounting for FTQC-specific overheads.
@@ Acknowledgements
We sincerely thank our shepherd and the anonymous reviewers for their feedback. We thank Karl Jansen and Stefan Kühn from the Center for Quantum Technology and Applications (CQTA)- Zeuthen for supporting this work by providing access to IBM quantum resources. We also thank Ahmed Darwish and Dmitry Lugovoy for their contributions to this work. Funded by the Bavarian State Ministry of Science and the Arts as part of the Munich Quantum Valley (MQV).